# Robust Blockwise Random Pivoting:

### v3

@@ 5 Interpolation Matrix Construction
In this section, we consider the construction of the interpolation matrix $\mathbf{W}$ after skeleton selection. In particular, we are interested in exploiting information from the $O\left(ndk\right)$ skeleton selection algorithms in Table 1, i.e. the matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ in Algorithms 1, 3, and 4, to construct $\mathbf{W}$ efficiently in $O(nk^{2})$ time.
We begin by noting the optimal interpolation matrix that solves the least square problem
@@
so that we can compute such optimal $\mathbf{W}$ exactly in $O\left(ndk+dk^{2}+nk^{2}\right)$ time via QR decomposition on $\mathbf{X}_{S}$. However, the dominant cost $O(ndk)$ requires additional passes through $\mathbf{X}$ and can be prohibitive after the skeleton selection stage.
5.1 Exactly ID-revealing Algorithms
With the output matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ from Algorithms 1 and 4, the corresponding optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be computed in $O(nk^{2})$ time following Algorithm 5.
@@ 5.1 Exactly ID-revealing Algorithms
1:(a) $\bm{\pi}\in\mathfrak{S}_{n}$ (such that $S=\bm{\pi}\left(1:k\right)$), (b) $\mathbf{L}\in\mathbb{R}^{n\times k}$,
2:(a) $\mathbf{W}\in\mathbb{R}^{n\times k}$
@@
4:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
@@
The sequential (Algorithm 1) and blockwise (Algorithm 4) random pivoting, as well as their greedy pivoting variations, are exactly ID-revealing (Definition 1.2).
@@
We first observe that the difference between sequential/blockwise random versus greedy pivoting lies only in the skeleton subset $S$. That is, assuming the same skeleton selection $S$, the resulting matrices $\mathbf{L}$ from random and greedy pivoting would be the same. Therefore, it suffices to show that Proposition 5.1 holds for random pivoting in Algorithms 1 and 4.
For both Algorithms 1 and 4, recall that $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$. We observe that $\mathbf{L}\mathbf{Q}^{(t)\top}$ provides a rank-$k$ approximation for $\mathbf{X}$: $\left\|\mathbf{X}-\mathbf{L}\mathbf{Q}^{(t)\top}\right\|_{F}^{2}=\left\|\mathbf{d}^{(t)}\right\|_{1}=\mathcal{E}_{\mathbf{X}}\left(S\right)<\tau\left\|\mathbf{X}\right\|_{F}^{2}$. We will show that $\mathbf{L}\mathbf{Q}^{(t)\top}=\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}$ in exact arithmetic.
@@
The computation of $\mathbf{L}_{2}\mathbf{L}_{1}^{-1}$ involves solving a small $k\times k$ system $(n-k)$ times, which takes $O(nk^{2})$ time in general. Overall, given $\mathbf{L}$ from Algorithms 1 and 4, we have that Algorithm 5 constructs the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}=\mathbf{L}\mathbf{L}_{1}^{-1}$.
@@
5.2 Approximately ID-revealing Algorithms
In contrast to Algorithms 1 and 4, $\mathbf{L}$ from the sketchy pivoting algorithms in Algorithm 3 is insufficient for computing the optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ exactly, intuitively due to the loss of information in sketching (i.e. randomized dimension reduction).
Proposition 5.4.
The sketchy pivoting algorithms (Algorithm 3) are approximately ID-revealing.
Proposition 5.4 follows directly from the observation that applying Algorithm 5 on $\mathbf{L}$ from Algorithm 3 (i.e. $\mathbf{W}=\mathbf{L}\mathbf{L}_{1}^{-1}$) solves a sketched version of Equation 8:
(9) $\min_{\mathbf{W}\in\mathbb{R}^{n\times k}}\left\|\mathbf{X}\widehat{\bm{\Omega}}-\mathbf{W}\mathbf{X}_{S}\widehat{\bm{\Omega}}\right\|_{F}^{2}\quad\Rightarrow\quad\mathbf{W}=\mathbf{L}\mathbf{L}_{1}^{-1}=\mathbf{X}\widehat{\bm{\Omega}}\left(\mathbf{X}_{S}\widehat{\bm{\Omega}}\right)^{\dagger},$
where $\widehat{\bm{\Omega}}=\bm{\Omega}\left(:,1:k\right)\in\mathbb{R}^{d\times k}$ consists of the first $k$ columns in the randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ drawn in Algorithm 3. Unfortunately, such sketched least square estimation [49, 42, 15] is known to be suboptimal as long as $l<d$ [40].
Specifically for the interpolation matrix construction, the sketched estimation with a sample size as small as $k$ (i.e. without oversampling) in Algorithm 3 leads to large interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$, as numerically illustrated in Section 6.2 (e.g. comparing $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ as SkCPQR/SkLUPP-ID in Figure 2 and $\mathcal{E}_{\mathbf{X}}\left(S\right)$ as SkCPQR/SkLUPP in Figure 1 (right)).
@@ 5.2 Approximately ID-revealing Algorithms
1:(a) $\bm{\pi}\in\mathfrak{S}_{n}$ ($S=\bm{\pi}\left(1:k\right)$), (b) Sample size $l\geq k$ (or pre-computed $\mathbf{Y}\in\mathbb{R}^{n\times l}$)
2:(a) $\mathbf{W}\in\mathbb{R}^{n\times k}$
3:(Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$)
4:($\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$)
@@
6:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
@@
As an intuitive but effective remedy, the oversampled sketchy ID (OSID) in Algorithm 6 generalizes Algorithm 5 by increasing the sample size beyond $k$ (i.e. oversampling), which is shown to considerably alleviate such suboptimality of Equation 9 in interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ (cf. Section 6.2, SkCPQR/SkLUPP-ID versus -OSID in Figures 2, 3, 4, 5, and 6). Concretely, with a moderate multiplicative oversampling $l=O(k)$ (e.g. $l=2k$ in Figures 2, 3, 4, 5, and 6 is usually sufficient in practice), Algorithm 6 can be expressed as
@@
Leveraging the existing theories for randomized linear embedding, [49, Theorem 9 and 23] implies that with Equation 10, the gap between the interpolation and skeletonization errors converges as $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)=O\left({k^{2}}/{l}\right)$ when $\bm{\Omega}$ is a sparse embedding and $l\gg k^{2}$, whereas such convergence can be further improved to $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)=O\left({k}/{l}\right)$ with Gaussian embeddings (e.g. by [49, Theorem 6]).
Asymptotically, Algorithm 6 takes $O(nlk+nk^{2}+lk^{2})=O(nk^{2}+k^{3})$ time. Compared to Algorithm 5 without oversampling, the dominant cost $O(nk^{2})$ increases proportionally to the oversampling factor $l/k$. Nevertheless, such additional cost is affordable in practice (cf. Figures 2, 3, 4, 5, and 6 (right)) thanks to the parallelizability of matrix-matrix multiplications [23].
@@ 6.1 Data Matrices
In the experiments, we compare the accuracy and efficiency of various ID algorithms on several types of data matrices, including the matrix forms of natural image datasets and synthetic random matrices with varied spectra, outlined as follows.
@@
    We draw a synthetic random data matrix $\mathbf{X}\in\mathbb{R}^{2000\times 500}$ from the adversarial GMM described in Example 4.1 with $k=100$ cluster centers, denoted as GMM.
@@
    Let Gaussian-exp be a random dense data matrix of size $1000\times 1000$ with exponential spectral decay. We construct Gaussian-exp from its SVD $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$ where $\mathbf{U},\mathbf{V}\in\mathbb{R}^{1000\times 1000}$ are random unitary matrices drawn from the Haar measure, and $\bm{\Sigma}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{1000}\right)$ where $\sigma_{i}=1$ for all $i\leq 100$ and $\sigma_{i}=\max\left\{0.8^{i-100},10^{-5}\right\}$ for all $i>100$.
@@
    Let SNN be a random sparse non-negative (SNN) matrix [44] of size $1000\times 1000$ such that

### v5

@@ 5 Interpolation Matrix Construction
In this section, we consider the construction of the interpolation matrix $\mathbf{W}$ after skeleton selection. In particular, we are interested in leveraging information from the skeleton selection processes to construct $\mathbf{W}$ efficiently.
Since the optimal interpolation matrix solves the following least-square problem:
@@
we can compute such optimal $\mathbf{W}$ exactly in $O\left(ndk+dk^{2}+nk^{2}\right)$ time via QR decomposition on $\mathbf{X}_{S}$. However, the dominant cost $O(ndk)$ requires additional passes through $\mathbf{X}$ and can be prohibitive after the skeleton selection stage.
5.1 ID-revealing Algorithms
With the output matrix $\mathbf{L}\in\mathbb{R}^{n\times k}$ from Algorithms 1 and 2, the corresponding optimal interpolation matrix $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be computed in $O(nk^{2})$ time following Algorithm 3.
@@ 5.1 ID-revealing Algorithms
2:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
@@
The sequential (Algorithm 1) and blockwise (Algorithm 2) random pivoting, as well as their greedy pivoting variations, are ID-revealing (Definition 1.1).
@@
We first observe that the difference between sequential/blockwise random v.s. greedy pivoting lies only in the skeleton subset $S$. That is, assuming the same skeleton selection $S$, the resulting matrices $\mathbf{L}$ from random and greedy pivoting would be the same. Therefore, it suffices to show that Proposition 5.1 holds for random pivoting in Algorithms 1 and 2.
For both Algorithms 1 and 2, recall that $\mathbf{L}\leftarrow\mathbf{L}^{(t)}$. We observe that $\mathbf{L}\mathbf{Q}^{(t)\top}$ provides a rank-$k$ approximation for $\mathbf{X}$: $\left\|\mathbf{X}-\mathbf{L}\mathbf{Q}^{(t)\top}\right\|_{F}^{2}=\left\|\mathbf{d}^{(t)}\right\|_{1}=\mathcal{E}_{\mathbf{X}}\left(S\right)<\tau\left\|\mathbf{X}\right\|_{F}^{2}$. We will show that $\mathbf{L}\mathbf{Q}^{(t)\top}=\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}$ in exact arithmetic.
@@
The computation of $\mathbf{L}_{2}\mathbf{L}_{1}^{-1}$ involves solving a small $k\times k$ system $(n-k)$ times, which takes $O(nk^{2})$ time in general. Overall, given $\mathbf{L}$ from Algorithms 1 and 2, we have that Algorithm 3 constructs the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}=\mathbf{L}\mathbf{L}_{1}^{-1}$.
@@
5.2 Oversampled Sketchy ID for Sketchy Pivoting
In contrast to random pivoting, sketchy pivoting [23, Algorithm 1] fails to be ID-revealing, intuitively due to the loss of information in the sketching process. Concretely, applying Algorithm 3 on sketchy pivoting leads to a sketched version of (6):
(7) $\min_{\mathbf{W}\in\mathbb{R}^{n\times k}}\left\|\mathbf{X}\widehat{\bm{\Omega}}-\mathbf{W}\mathbf{X}_{S}\widehat{\bm{\Omega}}\right\|_{F}^{2}\quad\Rightarrow\quad\mathbf{W}=\mathbf{X}\widehat{\bm{\Omega}}\left(\mathbf{X}_{S}\widehat{\bm{\Omega}}\right)^{\dagger},$
where $\widehat{\bm{\Omega}}=\bm{\Omega}\left(:,1:k\right)\in\mathbb{R}^{d\times k}$ consists of the first $k$ columns in the randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ drawn in sketchy pivoting. Unfortunately, such sketched least-square estimation [64, 59, 22] is known to be suboptimal as long as $l<d$ [56]. Specifically for interpolation matrix construction, (7) without oversampling leads to large interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\gg\mathcal{E}_{\mathbf{X}}\left(S\right)$, as numerically illustrated in Section 6.2 (cf. SkLUPP-ID in Figure 2 for $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ v.s. SkLUPP in Figure 1 for $\mathcal{E}_{\mathbf{X}}\left(S\right)$ (right)).
@@ 5.2 Oversampled Sketchy ID for Sketchy Pivoting
1:(Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ and compute $\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$)
@@
3:$\mathbf{W}\leftarrow\textbf{0}_{n\times k}$, $\mathbf{W}\left(\bm{\pi}\left(1:k\right),:\right)=\mathbf{I}_{k}$
@@
As an effective remedy, the oversampled sketchy ID (OSID) in Algorithm 4 generalizes Algorithm 3 by increasing the sample size beyond $k$ (i.e. oversampling), which is shown to considerably alleviate such suboptimality of (7) in interpolation error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ (cf. Section 6.2, SkLUPP-ID v.s. SkLUPP-OSID in Figures 2, 3, 4, 5, 6, and 7). Concretely, with a moderate multiplicative oversampling $l=\Theta(k)>k$ (e.g. $l=3k$ is usually sufficient in practice, see Section 6.2), Algorithm 4 can be expressed as
@@
Algorithm 4 takes $O(nlk+nk^{2}+lk^{2})=O(nk^{2})$ time. Compared to Algorithm 3 without oversampling, the dominant cost $O(nlk)$ increases proportionally to the oversampling factor $l/k$. Nevertheless, such additional cost is affordable in practice (cf. Figures 2, 3, 4, 5, 6, and 7) thanks to the hardware efficiency of matrix-matrix multiplications [34].
While sketchy pivoting without oversampling in (7) is not ID-revealing, the existing theories for sketch-and-solve (see e.g. [64, §2.5]) implies that OSID in (8) with Gaussian embedding and a multiplicative oversampling $l=\Theta(k)$ is $\gamma$-ID-revealing (recall Definition 1.1) for $\gamma=(1+O(\sqrt{k/l}))^{2}$ with a constant probability:
Proposition 5.4 (Sketchy pivoting + OSID is $\gamma$-ID-revealing).
Selecting a skeleton subset $S$ of size $k$ via sketchy pivoting [23, Algorithm 1] outputs (i) a sketched sample matrix $\mathbf{Y}=\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$ constructed by a Gaussian embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ with $l>k$ and i.i.d. entries from $\mathcal{N}(0,1/l)$, along with (ii) a permutation $\bm{\pi}\in\mathfrak{S}_{n}$ that characterizes the skeleton subset $S=\bm{\pi}\left(1:k\right)$. For any $\delta\in(0,1)$, when $l\gtrsim k+\log(1/\delta)$, with probability at least $1-\delta$, Algorithm 4 constructs an interpolation matrix $\mathbf{W}$ that satisfies $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\leqslant(1+O(\sqrt{\delta^{-1}k/l}))^{2}\mathcal{E}_{\mathbf{X}}\left(S\right)$ in $O(nkl)$ time.
Proof 5.5 (Proof of Proposition 5.4).
The proof of Proposition 5.4 is a direct generalization of [64, Theorem 23]. The key fact is that for a $(1/2,\delta/2,k)$-JLT [64, Definition 3] $\bm{\Omega}\in\mathbb{R}^{d\times l}$ with $(\Theta(\epsilon/\sqrt{k}),\delta/2)$-JL second moment property [64, Definition 12] for some $\epsilon>0$, (8) provides an interpolation matrix $\mathbf{W}$ that satisfies $\|\mathbf{W}\mathbf{X}_{S}-\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}\|_{F}\leqslant\epsilon\|\mathbf{X}-\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}\|_{F}$ (see e.g. the proofs of [64, Theorem 23] or [25, Lemma C.5]), and therefore, $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\leqslant(1+\epsilon)^{2}\mathcal{E}_{\mathbf{X}}\left(S\right)$. Then, it is sufficient to recall that a Gaussian embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$ with $l\gtrsim k+\log(1/\delta)$ is a $(1/2,\delta/2,k)$-Johnson-Lindenstrauss transform [64, Theorem 6] and satisfies the $(\Theta(\epsilon/\sqrt{k}),\delta/2)$-JL second moment property when $l\gtrsim\delta^{-1}k/\epsilon^{2}$.
Notice that Proposition 5.4 can be easily extended to various structured randomized linear embeddings like the sparse embedding in [11] and the LESS embedding in [18, 9] that offer better efficiency for constructing the sample matrix $\mathbf{Y}$.
@@ 6 Experiments
The experiments in Figures 1, 2, 3, 4, 5, 6, and 7 are conducted on a 14-core Apple M3 processor, while experiments in Table 3 are conducted on an Nvidia V100 GPU.
@@ 6.1 Data Matrices
We compare accuracy and efficiency of various ID algorithms on different types of data matrices, including natural data matrices from image datasets and integral operators, as well as synthetic matrices with varied spectra, outlined as follows.
@@
    We construct a synthetic data matrix $\mathbf{X}\in\mathbb{R}^{2000\times 500}$ based on the adversarial GMM example described in Example 4.1 with $k=100$ cluster centers, denoted as GMM.
@@
    Let Gaussian-exp be a random dense data matrix of size $1000\times 1000$ with exponential spectral decay. We construct Gaussian-exp from its SVD $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$ where $\mathbf{U},\mathbf{V}\in\mathbb{R}^{1000\times 1000}$ are random unitary matrices drawn from the Haar measure, and $\bm{\Sigma}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{1000}\right)$ where $\sigma_{i}=1$ for all $i\leqslant 100$ and $\sigma_{i}=\max\left\{0.8^{i-100},10^{-5}\right\}$ for all $i>100$.
@@
    SNN is a sparse non-negative (SNN) matrix [60] of size $1000\times 1000$ generated by
@@
  6. 6.
    Let Helmholtz be a matrix $\mathbf{X}$ induced by the Helmholtz kernel:
    $G(\mathbf{x},\mathbf{y})=\exp\left(i\kappa\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\right)/\left(4\pi\left\|\mathbf{x}-\mathbf{y}\right\|_{2}\right)\quad\text{with}\quad\kappa=5.5.$
    Consider a set of source points ${\mathcal{S}}$ consisting of $15\times 15\times 15$ Clenshaw-Curtis quadrature nodes in a 3D cube of size $2\times 2\times 2$ centered at the origin, along with a target set ${\mathcal{T}}$ containing $2000$ uniformly distributed points on the surface of a 3D sphere with radius $3$, also centered at the origin. We construct $\mathbf{X}$ of size $3375\times 2000$ by evaluating the Helmholtz kernel on ${\mathcal{S}}\times{\mathcal{T}}$ such that $\mathbf{X}_{ij}=G(\mathbf{x}_{i},\mathbf{y}_{j})$ for every $\mathbf{x}_{i}\in{\mathcal{S}}$ and $\mathbf{y}_{j}\in{\mathcal{T}}$.
It is worth highlighting that ID on natural data like MNIST, CIFAR-10, and Helmholtz is closely related to practical applications like data selection and fast direct solvers.
Remark 6.1 (ID for data selection).
ID on a training dataset like MNIST and CIFAR-10 can be viewed as an idealized data selection problem [1, 25]. For example, consider data selection for an overparametrized ridge regression problem without noise: given a full dataset $(\mathbf{X},\mathbf{y})\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$ satisfying $\mathbf{y}=\mathbf{X}\bm{\theta}_{*}$ for some unknown ground truth $\bm{\theta}_{*}\in\mathbb{S}^{d-1}$, we aim to approximate $\bm{\theta}_{*}$ via $\widehat{\bm{\theta}}=\mathop{\mathrm{argmin}}_{\bm{\theta}\in\mathbb{R}^{d}}\|\mathbf{X}_{S}\bm{\theta}-\mathbf{y}_{S}\|_{2}^{2}+\alpha\|\bm{\theta}\|_{2}^{2}$ with some $\alpha>0$ on a selected data subset $(\mathbf{X}_{S},\mathbf{y}_{S})\in\mathbb{R}^{k\times d}\times\mathbb{R}^{k}$ of size $k=\left|S\right|\leqslant d$. In this setting, the skeletonization error of ID, $\mathcal{E}_{\mathbf{X}}\left(S\right)$, is closely related to the generalization error of data selection:
$\text{as}\ \alpha\to 0,\quad\|\mathbf{X}(\widehat{\bm{\theta}}-\bm{\theta}_{*})\|_{2}^{2}=\|(\mathbf{X}\mathbf{X}_{S}^{\dagger}\mathbf{X}_{S}-\mathbf{X})\bm{\theta}_{*}\|_{2}^{2}\leqslant\mathcal{E}_{\mathbf{X}}\left(S\right).$
Beyond linear models like ridge regression, this connection can be extended to the more general finetuning problems for nonlinear models (like neural networks) by replacing $\mathbf{X}$ with the high-dimensional representations or model gradients evaluated at the data [25].
Remark 6.2 (ID for fast direct solvers).
In fast direct solvers for linear elliptic PDEs, ID is a widely used tool for compressing certain off-diagonal blocks of a rank-structured matrix [51, 32, 39, 40, 13, 53]. In this setting, both a high-quality skeleton selection and an accurate interpolation matrix construction are crucial for attaining a nearly minimum compression error. The ID experiments on the Helmholtz matrix are designed to model a compression step in the state-of-the-art fast direct solvers for Lippmann-Schwinger equation.